# Strategic resource allocation in memory encoding: An efficiency principle shaping language processing

### v1

@@ 3.2.1 All Types of Relations
In the analysis of the full dataset with all types of dependency relations, we indeed found a significant positive effect of antecedent surprisal for the majority of the languages, whereby more surprising antecedents are associated with longer dependency length. When dependency length $L$ is measured as intervening word counts, there is a robust positive effect in six languages, namely Danish, English, German, Italian, Russian, and Spanish. When $L$ is measured as intervening surprisal, in addition to these six languages there is Mandarin that also shows a significant positive effect. However, for Japanese, Korean and Turkish, contrary to our prediction, there is a negative antecedent surprisal effect, although this unexpected negative effect is less robust for the information-theoretic $L$.
@@
The result is quite mixed for the effect of antecedent frequency. When $L$ is measured as word counts, there is a negative effect of antecedent frequency on $L$ in six languages (robust for English, Italian, Russian, and Spanish, while partially supported in Korean and Turkish). The effect, however, is unexpectedly positive in German, Japanese, and Mandarin, and is inconclusive in Danish. When $L$ is measured as surprisal, the antecedent frequency effect unexpectedly turns out to be positive for more languages, namely Danish, English, German, Italian, Japanese and Mandarin (six languages). There is still a negative effect for Spanish and Turkish, although the effect is less robust compared to when $L$ is measured as word counts.
@@ 3.2.2 Subject Relations
For subject relations, when $L$ is measured as word counts, there is a positive effect of antecedent surprisal on $L$ as predicted in five languages (English, Italian, Russian, Spanish, and Mandarin), suggesting that more surprising antecedents are associated with longer $L$ in these languages. In contrast, contrary to our prediction, the effect is negative for German, Japanese, Korean, and Turkish, while Danish shows inconclusive result. When $L$ is measured as intervening surprisal, there are seven languages that exhibit a positive antecedent surprisal effect on $L$ (English, Italian, Russian, Spanish, Japanese, Korean, and Mandarin), representing the majority of languages in our dataset. However, the effect remains negative in German, and is inconclusive for Danish and Turkish.
@@
When $L$ is measured as word counts, there is a negative effect of antecedent frequency on $L$ in six languages (Danish, English, Italian, Russian, Spanish, and Korean), suggesting that more frequent antecedents lead to shorter $L$ in these languages. The effect is unexpectedly positive in German, and is inconclusive for Japanese, Mandarin, and Turkish. When $L$ is measured as surprisal, the result is similar to the orthographic $L$, except that the negative effect in Italian and Spanish becomes less robust and the effect in Korean becomes inconclusive.
@@ 3.2.3 Object Relations
The result for object relations shows a mixed pattern. When $L$ is measured as word counts, no positive effect of antecedent surprisal on $L$ is observed in any language. Instead, there is a negative effect in German, Russian, Spanish, Korean, and Mandarin, and the result is inconclusive for the rest of the languages. However, the pattern is reversed for the information-theoretic $L$, with a positive antecedent surprisal effect being observed in seven languages (Danish, English, Italian, Russian, Spanish, Japanese, Turkish). Only in Korean did we observe a negative effect, and the result is inconclusive in German and Mandarin.
@@
When $L$ is measured as word counts, we only found a robust antecedent frequency effect on $L$ in four languages, two negative (Italian and Spanish) and two positive (English and German). The result for the rest of the languages is inconclusive. When $L$ is measured as intervening surprisal, there are more languages that show an unexpectedly positive effect (Danish, English, German, Italian, Korean, and Mandarin). Only in Spanish did we observe a negative antecedent frequency effect. The result for Russian, Japanese and Turkish remains inconclusive.
@@ 3.3 Discussion
In this cross-linguistic corpus study, we indeed found evidence for a positive effect of antecedent surprisal on dependency length $L$, especially when $L$ is measured as intervening surprisal (bottom left panel in Figure 5), which, as mentioned above, should better capture the variability of information load across word types. With this information-theoretic dependency length measure, the majority of languages (seven out of ten) exhibit a positive effect of antecedent surprisal on $L$, and this effect still holds when we zoom into the subset that only includes subject or object relations. Overall, as predicted, this pattern indicates that more surprising antecedents are associated with longer dependency length, suggesting that the pressure to minimize dependency length is relaxed when the antecedent is of higher surprisal. Consistent with our hypothesis of strategic resource allocation, the result supports that novel and unexpected linguistic units can tolerate longer dependency length before its retrieval site, possibly because unexpected information is prioritized for working memory resources during encoding, and is more resistant to memory decay and interference.
However, there are two caveats worth noting. First, there is considerable cross-linguistic variability in our result, and the antecedent surprisal effect mostly exists within the Indo-European and head-initial languages in our analysis. Second, although the analysis on the full dataset with all types of dependencies reveals a general trend for a positive antecedent surprisal effect, the result is much more consistent within subject relations across different dependency length measures than within object relations. In object relations, the expected effect is reversed with the orthographic dependency length.
It is also worth noting that the positive antecedent surprisal effect on $L$ cannot be reduced to a pure frequency effect. That is, there is still a significant effect of antecedent surprisal even though antecedent frequency has been included in the regression models as a control variable. Moreover, compared to antecedent surprisal, the effect of antecedent frequency on $L$ is much less consistent, and better aligns with our prediction with the orthographic $L$. That is, more frequent antecedents are associated with shorter $L$ mostly when $L$ is measured as word counts instead of intervening surprisal. Taken together, the antecedent surprisal effect we found on dependency length is irreducible to a pure frequency effect, and the antecedent frequency effect by itself is less consistent.
In the end, to what extent does written corpus text approximate language production? Compared to spoken language, written language typically allows “speakers” more time to think, reducing much of the cognitive load involved in production, and the communicative goal is more geared towards listeners’ need. That being said, speaker-oriented cognitive constraints, such as memory capacity, may play a less prominent role, and the need for strategic memory allocation may be diminished in written language production. Therefore, the effect observed in the current analysis using written text can be viewed as an upper bound, and we expect the effect of strategic memory allocation to be stronger when using spoken language corpora.
@@ 4 Study 2a: Comprehension Side (Self-Paced Reading)
In this second study, we investigate whether the effect of strategic resource allocation also holds from the comprehension side. In particular, we examine to what extent the dependency locality effect observed in previous studies (Gibson, 1998, 2000; Grodner & Gibson, 2005; Bartek et al., 2011) can be modulated by the surprisal of the antecedent. First, we expect a baseline dependency locality effect, where the processing difficulty at the retrieval site, manifested as reading time, is expected to increase as the dependency length gets longer. Second, according to our hypothesis, more surprising antecedents are more capable of tolerating stronger memory interference. That is, the longer dependency length does not create too much additional processing difficulty for the retrieval of surprising antecedents, resulting in a reduced locality effect. Consider the example below in (6):
  1. (6)
    1. a.
      The evil monster{ $\text{The evil {monster}}\begin{Bmatrix}\text{ in the tower }\\ \text{ who stayed in the tower }\\ \text{ who stayed in the tower near the castle }\end{Bmatrix}\text{\text@underline{approached}...}$
    2. b.
      The cute monster{ $\text{The cute {monster}}\begin{Bmatrix}\text{ in the tower }\\ \text{ who stayed in the tower }\\ \text{ who stayed in the tower near the castle }\end{Bmatrix}\text{\text@underline{approached}...}$
As explained in Study 1, “the cute monster” in (6b) is more prioritized with encoding resources for its lower predictability, making it capable of tolerating longer dependency length. Therefore, when there is longer dependency distance, the processing difficulty at the retrieval site (i.e., the main verb “approached” in this example) should increase at slower rate in (6b) than in (6a), since the additional intervening material induces lower level of interference for the more surprising antecedent in (6b). As a result, on top of the baseline effect on retrieval difficulty, we expect a negative interaction between dependency length $L$ and antecedent surprisal at the retrieval site.
@@ 4.1.1 Data
The data we used in Study 2a is taken from the Natural Stories Corpus (NSC) (Futrell et al., 2021). The text of the corpus is in English, and contains 10,245 lexical words in 485 sentences, taken from 10 stories with around 1000 words each. The corpus includes dependency annotation. The reading time (RT) data is collected from 181 native English speakers, using the self-paced reading task (SPR). The original corpus already excluded participants with low comprehension accuracy, as well as the reading times either faster than 100ms or slower than 3000ms. Therefore, we did not perform additional exclusion of reading time data in the current study. We generated the surprisal estimates for each word from mGPT. [^10] GPT-3 is no longer accessible from OpenAI since January 4th, 2024. Therefore, we only used the surprisal estimates from mGPT for the analyses in Study 2a and 2b. The NSC corpus comes with dependency annotation in UD style.
@@ 4.1.2 Data Transformation, Exclusion, and Analysis
As in Study 1, we analyzed three RT datasets as well here in the current study, namely the full dataset with all types of dependencies, a subset with subject relations only, and a subset with object relations only. The sample size is summarized in Table 2. We ran linear mixed-effect models on the log-transformed RTs at the right codependent of each dependency, which is considered the retrieval site for the antecedent. The critical effect is the interaction between dependency length $L$ and antecedent surprisal, with maximal converging random intercept and random slopes by participant. For the analysis of the full dataset, we also included maximal converging random effects by dependency type.
@@
The four control variables applied in Study 1 (namely, sentence position, antecedent position, sentence length, and antecedent frequency) are also included here in Study 2a. Besides these, we included several additional control variables that are often considered highly relevant for reading time measures. First, we included word length, word surprisal and word frequency of the right codependent itself. Second, to control the spillover effect often found in reading studies, we included word surprisal and word frequency of the two previous words before the right codependent. As in Study 1, frequency measures are in log scale, generated from Speer (2022). All variables are $z$-scaled. The transformation and exclusion of the dependency data follow the same procedure as in Study 1.
@@ 4.2 Results
Figure 6: Study 2a reading time (RT) result. Panel A: raw RTs at the right codependent as a function of dependency length $L$ modulated by antecedent surprisal, which is binned into tertiles for visualization. Panel B: result of regression models with log-transformed RTs; coefficient estimates with 95% confidence interval for the effects of $L$, antecedent surprisal, and the interaction between the two. Significance levels: *(p$<$0.05), **(p$<$0.01), ***(p$<$0.001)
@@
In the analysis of the full dataset with all types of dependency relations, we first found a baseline locality effect, where dependency length $L$ has a positive main effect on RTs at the retrieval site (i.e., the right codependent), suggesting that longer distance between codependents makes it more difficult to retrieve the antecedent at the right codependent. This baseline locality effect holds for both measures of $L$. However, there is no evidence for an antecedent surprisal main effect. In the end, importantly, we found a negative $L$$\times$ antecedent surprisal two-way interaction. Consistent with our prediction, this negative interaction suggests that the locality effect on the RT of right codependents is reduced when the antecedent is more surprising. This two-way interaction, however, only holds when $L$ is measured as surprisal, and no interaction is found when $L$ is measured as word counts.
@@
First, for both measures of dependency length $L$, a baseline locality main effect of $L$ was found for subject relations, where longer $L$ leads to longer RT at the right codependent. Second, as in the analysis of the full dataset above, there is no antecedent surprisal main effect with either measure of $L$. In the end, we found a critical negative $L$$\times$ antecedent surprisal two-way interaction with both $L$ measures. As predicted, this negative interaction is indicative of a reduced locality effect for more surprising antecedents.
@@
The result of object relations has a relatively complex pattern. First, unlike the previous two analyses, there is an unexpected negative, not positive, main effect of $L$ on the RT at retrieval site. This negative main effect reaches significance when $L$ is measured as surprisal, and still numerically holds when $L$ is measured as word counts. Instead of a baseline locality effect, this negative $L$ effect suggests an anti-locality effect, where more intervening material between the two codependents leads to faster RT at the retrieval site. Moreover, as seen in Figure 6A (right column), this anti-locality effect is mostly driven by antecedents of low surprisal. Second, there is a negative antecedent surprisal main effect for both $L$ measures, whereby more surprising antecedents induce faster RT at the retrieval site. In the end, unlike the negative $L$$\times$ antecedent surprisal interaction observed in the previous two analyses, object relations exhibit a positive $L$$\times$ antecedent surprisal interaction, which is significant for both $L$ measures. However, since the main effect of $L$ is negative in the first place, instead of an enhanced locality effect, this positive interaction actually suggests a reduced anti-locality effect for more surprising antecedents. It is not yet entirely clear to us why there is an anti-locality effect in the first place exclusively for object relations, but the result pattern in object relations seems to point to a potential trade-off between the direction of the $L$ main effect and the direction of the $L$$\times$ antecedent surprisal interaction, which we will discuss below in Section 4.3.
@@ 4.3 Discussion
To sum up, in Study 2a, through the analysis of the RT data in the Natural Stories Corpus, we first replicated the baseline locality effect in the analysis of all types of dependency relations and in the one of subject relations only, suggesting that the non-local retrieval of the antecedent at the right codependent becomes more difficult when there is more intervening material. However, we also found an anti-locality effect in object relations, which suggests that more intervening material actually facilitates the establishment of a non-local object relation. Second, only in object relations did we observe a negative main effect of antecedent surprisal on the RT at the right codependent, suggesting that more surprising antecedents are easier to retrieve in object relations. This negative effect of antecedent surprisal is consistent with the findings in Hofmeister (2011), which observes that semantically more complex noun phrases are easier to be retrieved. However, more future work is needed to investigate why this effect only emerges in object relations. In the end, the most important finding of this Study 2a is the interaction between dependency length and antecedent surprisal. Specifically, in the analysis of the full dataset and the one of subject relations, we found a reduced locality effect for more surprising antecedents. This reduced locality effect is consistent with the prediction of strategic resource allocation, in that more surprising antecedents are prioritized for working memory resources and are encoded with more robust representation against memory interference and decay.
@@
It is also worth noting that there seems to be a potential trade-off between the direction of the $L$ main effect and the direction of the $L$$\times$ antecedent surprisal interaction. On the one hand, when the main effect of $L$ is positive, as in the first two analyses, there is a negative $L$$\times$ antecedent surprisal interaction, suggesting a reduced locality effect for more surprising antecedents. On the other hand, when the main effect of $L$ is negative to start with, as in the object relations, the interaction becomes positive, indicating that the anti-locality effect is reduced for more surprising antecedents. The reduced locality effect is straightforward, as predicted by our hypothesis. But why is there a reduced anti-locality effect? In fact, the reduced anti-locality effect for more surprising antecedents can be consistent with our strategic resource allocation as well. According to experience-based processing theories, the anti-locality effect can be viewed as a facilitation effect on the prediction of the right codependent. That is, more intervening material may provide more information about the identity of the word at the right codependent, helping the comprehender to make better predictions (Levy, 2008a), thus canceling out the burden created by memory interference. However, for more surprising antecedents, if their representation is more enhanced due to strategic resource allocation, it is possible that comprehenders can already rely on the their memory of the antecedent to predict the right codependent. As a result, the intervening material may no longer provide too much additional help to make predictions. This reduced facilitation from the intervening material for more surprising antecedent, therefore, may manifest itself as a reduced anti-locality effect.

### v2

@@ 4.2.1 All Types of Relations
In the analysis of the full dataset with all types of dependency relations, we indeed found a significant positive effect of antecedent surprisal for six out of ten languages, whereby more surprising antecedents are associated with longer dependency length. Specifically, for both measures of dependency length $L$, there is a positive effect in Danish, English, German, Italian, Russian, and Spanish. However, for Japanese, Korean, Mandarin and Turkish, contrary to our prediction, there is a negative antecedent surprisal effect.
@@
The result is quite mixed for the effect of antecedent frequency. When $L$ is measured as intervening word counts, there is a negative effect of antecedent frequency on $L$ in six languages (robust for English, Italian, Russian, and Spanish, while partially supported in Korean and Turkish). The effect, however, is unexpectedly positive in German, Japanese, and Mandarin, and is inconclusive in Danish. When $L$ is measured as surprisal, the antecedent frequency effect unexpectedly turns out to be positive for more languages, namely Danish, English, German, Japanese and Mandarin. There is still a negative effect for Russian, Spanish, Korean and Turkish.
@@ 4.2.2 Subject Relations
For subject relations, when $L$ is measured as word counts, there is a positive effect of antecedent surprisal on $L$ as predicted in five languages (English, Italian, Russian, Spanish, and Mandarin), suggesting that more surprising antecedents are associated with longer $L$ in these languages. However, contrary to our prediction, the effect is negative for German, Japanese, Korean, and Turkish, while Danish shows inconclusive result. Similar pattern was observed when $L$ is measured as intervening surprisal, except for Mandarin whose effect becomes inconclusive.
@@
The effect of antecedent frequency is the same for both measures of $L$. That is, there is a negative effect of antecedent frequency on $L$ in six languages (Danish, English, Italian, Russian, Spanish, and Korean), suggesting that more frequent antecedents lead to shorter $L$ in these languages. The effect is unexpectedly positive in German, and is inconclusive for Japanese, Mandarin, and Turkish.
@@ 4.2.3 Object Relations
Surprisingly, for object relations, there is no positive effect of antecedent surprisal on $L$ in any language when $L$ is measured as word counts. Instead, there is a negative effect in German, Russian, Spanish, Korean, and Mandarin, and the result is inconclusive for the rest of the languages. Similar pattern was observed when $L$ is measured as surprisal, except that in Italian a positive effect is partially supported, and that the originally negative effect with orthographic $L$ in Russian and Spanish becomes less robust.
@@
There is also a mixed picture for the effect of antecedent frequency in object relations. When $L$ is measured as word counts, we only found a robust antecedent frequency effect on $L$ in four languages, two negative (Italian and Spanish) and two positive (English and German). The result for the rest of the languages is inconclusive. When $L$ is measured as intervening surprisal, there are three languages that show an unexpectedly positive effect (English, German, and Mandarin). Only in Spanish did we observe a negative antecedent frequency effect. The result for the rest of the languages remains inconclusive.
@@ 4.3 Discussion
In this cross-linguistic corpus study, we indeed found emerging evidence for a positive effect of antecedent surprisal on dependency length $L$, with both measures of $L$ showing similar patterns. This effect still holds when we zoom into the subset that only includes subject or object relations. Overall, in many languages (especially Indo-Europeans), as predicted, this pattern indicates that more surprising antecedents are associated with longer dependency length, suggesting that the pressure to minimize dependency length is relaxed when the antecedent is of higher surprisal. Consistent with our hypothesis of strategic resource allocation, the result supports that novel and unexpected linguistic units can tolerate longer dependency length before its retrieval site, possibly because unexpected information is prioritized for working memory resources during encoding, and is more resistant to memory decay and interference.
However, there are two caveats worth noting. First, there is considerable cross-linguistic variability in our result, and the antecedent surprisal effect mostly exists within Indo-European and head-initial languages in our analysis. Second, although the analysis on the full dataset with all types of dependencies reveals a general trend for a positive antecedent surprisal effect, the result is much more consistent within subject relations. In object relations, the expected effect is reversed for most languages.
It is also worth noting that the positive antecedent surprisal effect on $L$ cannot be reduced to a pure frequency effect. That is, there is still a significant effect of antecedent surprisal even though antecedent frequency has been included in the regression models as a control variable. Moreover, compared to antecedent surprisal, the effect of antecedent frequency on $L$ is less consistent.
In the end, to what extent does written corpus text approximate language production? Compared to spoken language, written language typically allows “speakers” more time to think, reducing much of the cognitive load involved in production, and the communicative goal is more geared towards listeners’ need. That being said, speaker-oriented cognitive constraints, such as memory capacity, may play a less prominent role, and the need for strategic memory allocation may be diminished in written language production. Therefore, the effect observed in the current analysis using written text can be viewed as an lower bound, and we expect the effect of strategic memory allocation to be stronger when using spoken language corpora.
@@ 5 Study 2a: Comprehension Side (Self-Paced Reading)
In this second study, we investigate whether the effect of strategic resource allocation also holds from the comprehension side. In particular, we examine to what extent the dependency locality effect observed in previous comprehension studies (Gibson, 1998, 2000; Grodner & Gibson, 2005; Bartek et al., 2011) can be modulated by the surprisal of the antecedent, as illustrated in Figure 8. First, we expect a baseline dependency locality effect, where the processing difficulty at the retrieval site, manifested as reading time, is expected to increase as the dependency length gets longer. Second, according to our hypothesis, more surprising antecedents are more capable of tolerating stronger memory interference. That is, the longer dependency length does not create too much additional processing difficulty for the retrieval of surprising antecedents, resulting in a reduced locality effect.
Figure 8: Empirical prediction of strategic resource allocation in comprehension. Longer dependency length leads to higher processing difficulty at the retrieval site, and this retrieval difficulty increases more slowly for high-surprisal antecedents.
Consider the example in Figure 8. As explained in Study 1, the more surprising “cute monster” in Figure 8 (bottom) is more prioritized with encoding resources for its lower predictability, making it capable of tolerating longer dependency length. Therefore, when there is longer dependency distance, the processing difficulty at the retrieval site (i.e., the main verb “approached” in this example) should increase at slower rate for high surprisal antecedents than for low surprisal ones (Figure 8; top), since the additional intervening material induces lower level of interference for more surprising antecedents. As a result, on top of the baseline dependency length effect on retrieval difficulty, we expect a negative interaction between dependency length $L$ and antecedent surprisal at the retrieval site.
@@ 5.1.1 Data
The data we used in Study 2a is taken from the Natural Stories Corpus (NSC) (Futrell et al., 2021). The text of the corpus is in English, and contains 10,245 lexical words in 485 sentences, taken from 10 stories with around 1000 words each. The reading time (RT) data was collected from 181 native English speakers, using the self-paced reading task (SPR). The original corpus already excluded participants with low comprehension accuracy, as well as the reading times either faster than 100ms or slower than 3000ms. Therefore, we did not perform additional exclusion of reading time data in the current study. We generated the surprisal estimates for each word from mGPT. [^16] GPT-3 is no longer accessible from OpenAI since January 4th, 2024. Therefore, we only used the surprisal estimates from mGPT for the analyses in Study 2a and 2b. The NSC corpus comes with dependency annotation in UD style.
@@ 5.1.2 Data Transformation, Exclusion, and Analysis
As in Study 1, we analyzed three RT datasets as well here in the current study, namely the full dataset with all types of dependencies, a subset with subject relations only, and a subset with object relations only. The sample size is summarized in Table 2. We ran linear mixed-effect models on the log-transformed RTs of two regions, the critical region and the spillover region. The critical region is the right codependent of each dependency, which is considered the retrieval site for the antecedent. The spillover region is the word that goes immediately after the critical region. The critical effect is the interaction between dependency length $L$ and antecedent surprisal, with maximal converging random intercept and random slopes by participant. For the analysis of the full dataset, we also included maximal converging random effects by dependency type.
@@
The control variables applied in Study 1 (namely, sentence position, antecedent position, sentence length, and antecedent frequency) are also included here in Study 2a. Besides these, we included several additional control variables that are often considered highly relevant for reading time measures. First, we included word length, word surprisal and word frequency of the right codependent itself. Second, to control the spillover effect often found in reading studies, we included word surprisal and word frequency of the two previous words before the right codependent. As in Study 1, frequency measures are in log scale, generated from Speer (2022). All variables are $z$-scaled. The transformation and exclusion of the dependency data follow the same procedure as in Study 1.
@@ 5.2 Results
Figure 9: Study 2a reading time (RT) result. Panel A: raw RTs at the right codependent and its spillover region as a function of dependency length $L$ modulated by antecedent surprisal, which is binned into tertiles for visualization. Panel B: result of regression models with log-transformed RTs; coefficient estimates with 95% confidence interval for the effects of $L$, antecedent surprisal, and the interaction between the two. Significance levels: *(p<<0.05), **(p<<0.01), ***(p<<0.001)
@@ 5.2.1 Critical Region
In the analysis of the full dataset with all types of dependency relations, we found a baseline locality effect where dependency length $L$ has a positive main effect on RTs at the retrieval site (i.e., the right codependent), suggesting that longer distance between codependents makes it more difficult to retrieve the antecedent at the right codependent. This main effect of $L$ is only significant when $L$ is measured as intervening surprisal. However, there is no significant main effect of antecedent surprisal. Importantly, we found a negative $L$ $\times$ antecedent surprisal two-way interaction for both $L$ measures. Consistent with our prediction, this negative interaction suggests that the locality effect on the RT of right codependents is reduced when the antecedent is more surprising.
@@
For subject relation, although the main effect of $L$ is numerically positive in the critical region, it is not significant with any measures of $L$. There is no antecedent surprisal main effect, either. However, there is indeed a significant negative $L$ $\times$ antecedent surprisal two-way interaction for both $L$ measures, suggesting a reduced locality effect for high surprisal antecedents.
@@
Again, for object relations, there is no main effect of $L$ or antecedent surprisal for any measures of $L$ in the critical region. It is also worth noting that the $L$ main effect is numerically negative, pointing to an anti-locality effect, although this effect is not statistically significant. Surprisingly, there is a positive $L$ $\times$ antecedent surprisal two-way interaction, although this effect is only significant when $L$ is measured as word counts.
5.2.2 Spillover Region
All types of relations.
In the spillover region, we first found a baseline locality effect, where dependency length $L$ leads to longer RT. This baseline locality effect holds for both measures of $L$. However, similar to the critical region, there is no evidence for an antecedent surprisal main effect. In the end, again similar to the critical region, we found a negative $L$ $\times$ antecedent surprisal two-way interaction, suggesting that the locality effect is reduced when the antecedent is more surprising. This two-way interaction, however, only holds when $L$ is measured as surprisal.
Subject relations.
First, unlike the critical region, a baseline locality main effect of $L$ was found for subject relations in the critical region, where longer $L$ leads to longer RT at the right codependent. This $L$ main effect is significant for both measures of dependency length $L$. Second, as in the critical region, there is no antecedent surprisal main effect with either measure of $L$ in the spillover region. In the end, we found a critical negative $L$ $\times$ antecedent surprisal two-way interaction with both $L$ measures. As predicted, this negative interaction is indicative of a reduced locality effect for more surprising antecedents.
Object relations.
The result of object relations has a relatively complex pattern. First, there is an unexpected negative, not positive, main effect of $L$ on the RT at retrieval site. This negative main effect reaches significance when $L$ is measured as surprisal, and still numerically holds when $L$ is measured as word counts. Instead of a baseline locality effect, this negative $L$ effect suggests an anti-locality effect, where more intervening material between the two codependents leads to faster RT at the retrieval site. Moreover, as seen in Figure 9A (right column), this anti-locality effect is mostly driven by antecedents of low surprisal. Second, there is a negative antecedent surprisal main effect for both $L$ measures, whereby more surprising antecedents induce faster RT at the retrieval site. In the end, unlike the negative $L$ $\times$ antecedent surprisal interaction observed in the previous two analyses, object relations exhibit a positive $L$ $\times$ antecedent surprisal interaction, which is significant for both $L$ measures. However, since the main effect of $L$ is negative in the first place, instead of an enhanced locality effect, this positive interaction actually suggests a reduced anti-locality effect for more surprising antecedents. It is not yet entirely clear to us why there is an anti-locality effect in the first place exclusively for object relations, but the result pattern in object relations seems to point to a potential trade-off between the direction of the $L$ main effect and the direction of the $L$ $\times$ antecedent surprisal interaction, which we will discuss below.
@@ 5.3 Discussion
To sum up, in Study 2a, through the analysis of the RT data in the Natural Stories Corpus, we first replicated the baseline locality effect in the analysis of all types of dependency relations and subject relations, especially in the spillover region. This suggests that the non-local retrieval of the antecedent at the right codependent becomes more difficult when there is more intervening material. However, we also found an anti-locality effect in object relations in the spillover region, which suggests that more intervening material actually facilitates the establishment of a non-local object relation. Second, only in object relations did we observe a negative main effect of antecedent surprisal on the RT at the right codependent, but this effect only emerges in the spillover region. This negative antecedent surprisal main effect suggests that more surprising antecedents are easier to retrieve in object relations, which is consistent with the findings in Hofmeister (2011), where semantically more complex noun phrases are easier to be retrieved. However, more future work is needed to investigate why this effect only emerges in object relations. In the end, the most important finding of this Study 2a is the interaction between dependency length and antecedent surprisal, both in the critical and the spillover regions. Specifically, in the analysis of the full dataset and the one of subject relations, we found a reduced locality effect for more surprising antecedents. This reduced locality effect is consistent with the prediction of strategic resource allocation, in that more surprising antecedents are prioritized for working memory resources and are encoded with more robust representation against memory interference and decay.
@@
It is also worth noting that there seems to be a potential trade-off between the direction of the $L$ main effect and the direction of the $L$ $\times$ antecedent surprisal interaction. On the one hand, when the main effect of $L$ is positive, there is a negative $L$ $\times$ antecedent surprisal interaction, suggesting a reduced locality effect for more surprising antecedents. On the other hand, when the main effect of $L$ is negative to start with, as in the object relations, the interaction becomes positive, indicating that the anti-locality effect is reduced for more surprising antecedents. The reduced locality effect is straightforward, as predicted by our hypothesis. But why is there a reduced anti-locality effect? In fact, the reduced anti-locality effect for more surprising antecedents can be consistent with our strategic resource allocation as well. According to experience-based processing theories, the anti-locality effect can be viewed as a facilitation effect on the prediction of the right codependent. That is, more intervening material may provide more information about the identity of the word at the right codependent, helping the comprehender to make better predictions (Levy, 2008a), canceling out the burden created by memory interference. However, for more surprising antecedents, if their representation is more enhanced due to strategic resource allocation, it is possible that comprehenders can already rely on the their memory of the antecedent to predict the right codependent. As a result, the intervening material may no longer provide too much additional help to make predictions. This reduced facilitation from the intervening material for more surprising antecedent, therefore, may manifest itself as a reduced anti-locality effect.